OpenStudy (anonymous):

Your local cable television company offers two plans: basic service with one movie channel for $35 per month, or basic service with two movie channels for $45 per month. What is the charge for the basic service and the charge for each movie channel?

OpenStudy (anonymous):

35 then 10 each?/

OpenStudy (anonymous):

B+2M=45 <---basic plus 2 movie channels is $45 B+ M=35 <---basic plus 1 movie channel is $35 -------- <---subtract the second equation from the first M=10 <--- (because B-B=0, 2M-M=M, and 45-35=10) So if M=10, B+M=B+10 If B+10=35, what is B?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

i didnt catch that nice 1